What is the the energy (in Joules per photon) of ultraviolet radiation that has a wavelength of 1.667e-7 meters? The speed of light is 3.00x10 8 m/s and Planck's constant is 6.626x10 -34 J*s.

Key Concept:Energy of a photon is related to the frequency, ν, and wavelength, λ. E = h*ν = hc/λ where c is the speed of light.
1
Expert's answer
2020-12-21T03:30:32-0500

E= hc/λ

E = 6.626 ×10-34 × 3 × 108 / (1.667 ×10-7)

E = 1.19 × 10-18J